OpenStudy (anonymous):

what is the solution to the rational equation (3/2)-(1/(x-4))=(-2/(2x-8))

OpenStudy (anonymous):

a.) no solution b.) -4 c.) 2 d.) 4

hero (hero):

\(\frac{3}{2} - \frac{1}{x-4} = \frac{-2}{2x-8}\) Hint: Multiply the second fraction 1/(x-4) by 2/2

OpenStudy (anonymous):

2/2x-8

hero (hero):

Yes, and now you have \(\frac{3}{2} - \frac{2}{2x-8} = \frac{-2}{2x-8}\) What happens when you add \(\frac{2}{2x-8}\) to both sides?

OpenStudy (anonymous):

it cancels

hero (hero):

Well actually, you end up with \(\frac{3}{2} = -\frac{2}{2x-8} + \frac{2}{2x-8}\)

hero (hero):

Which basically becomes \(\frac{3}{2} = 0\) Which we know is not true, therefore, no solution

hero (hero):

Is there something you wanted to say?

hero (hero):

@millitarygirl96